# 4Université de Strasbourg, CNRS UMR 7550, Observatoire astronomique de Strasbourg, 67000 Strasbourg, France 5DIO, Observatoire de Paris, CNRS, PSL, 75014, Paris, France 6IDRIS, CNRS, F-91403 Orsay, France

### v1

@@ 5 Discussion
This second edition of the SKAO SDCs improves the challenge format in several aspects compared to the first one. The simulated cube includes more physical and instrumental effects, improving its realism, which is also a strength of the subsequent SDC editions (see the SDC3a, Bonaldi et al., 2025). The LDEV cube is a proper representation of the MAIN cube, allowing for training supervised ML methods without requiring representativity corrections like we implemented in Paper I. Still, the use of a similar scoring method to the SDC1 results in the same bias and limits, such as strong coupling between detection and characterization performance evaluation, or the absence of a confidence score weighting like in most computer-vision challenges. We also acknowledge that using simulated data provides an unrealistic advantage for training supervised ML methods compared to real observations. All sources are known and perfectly characterized, allowing for the definition of unambiguous targets, which would not be the case with observational labeled catalogs.
@@
The results obtained with our YOLO-CIANNA 3D method, combined with the fact that the two top-scoring teams from the SDC2 utilized 3D CNNs, demonstrate that this type of structure offers significant advantages for processing noisy hyperspectral data. However, as discussed in Paper I, this requires building dedicated approaches that account for the specificities of astronomical data deeply in their design, which was facilitated by the low-level access granted by our CIANNA framework. Nevertheless, several aspects of our method could be improved. In terms of backbone architecture, CIANNA V-1.0 does not implement residual and skip connection layers (He et al., 2016), which limit the maximum network depth. Such layers have been added to the experimental build of CIANNA, which is currently only provided upon specific request but is expected to be publicly released on a short timescale. Unlike some versions of YOLO (Redmon & Farhadi, 2018), we did not split predictions across scales, which is also related to current architectural limitations. In addition, having a fixed number of detection units on a finite-sized grid constrains both the positioning resolution and the density of detectable sources. This specific limit could likely be overcome with the addition of attention layers (Vaswani et al., 2017; Carion et al., 2020; Zhang et al., 2021) and some changes in the output layer definition. We also consider the possibility of pre-training our backbone using the Denoising Diffusion Probabilistic Model (DDPM, Ho et al., 2020) formalism. For example, we could have pre-trained our model over the MAIN cube without requiring any source targets. This would likely shorten the initial loss plateau, resulting in more stable training and possibly better results. Finally, we could inject context information, such as the instrumental setup, at multiple stages of the backbone to guide the source extraction (Lu et al., 2019). This would enable us to train generic detection models compatible with multiple setups or instruments.
@@
The natural next step after achieving state-of-the-art results on simulated data is to generalize our method application to observational data from SKA precursors such as the LADUMA (Looking at the Distant Universe with the MeerKAT Array, Blyth et al., 2016) and WALLABY (Widefield ASKAP L-band Legacy All-sky Blind surveY, Koribalski et al., 2020) surveys. Our method could also be generalized to detect line emission in hyperspectral surveys covering a different frequency domain, such as MUSE. However, it requires a large set of labeled examples to build the training and validation sets. Although simulations can be used to generate large example datasets for pre-training highly parametric statistical models, they often lose accuracy when generalized to real observations. A complementary training phase on real labeled data is almost always necessary, and the calibration of the inference pipeline must necessarily be done on a labeled validation dataset. For this, we can combine labels from classical detection methods, crowdsourced science, or existing catalogs from other instruments. The supervised nature of our method implies that the trained detector will likely inherit uncertainties and biases from the methods used to define the targets. Combining independent labeling methods might help in this regard. In all cases, the bootstrap training approach can be employed to identify candidate detection that can then be refined using different labeling approaches or new observational programs, iteratively improving detection sensitivity. Preparatory works indicate that direct porting of our SDC2-trained models generalizes well over preliminary LADUMA data after a simple rescaling, producing candidate detection that is compatible with catalogs obtained with SOFIA. Specific retraining for this survey is currently being explored.
We conclude with an update of the numerical environmental footprint estimated in Paper I, which already accounts for most of the impact related to the method development and testing for 2D and 3D datasets. For this SDC2 study, we estimate the additional compute time at 4000 GPU hours, which includes the 1000 hours granted to MINERVA on Jean-Zay during the original challenge. Using the same estimates of 0.5 kW of average power consumption when training with a single GPU, and France’s carbon intensity estimated by the ADEME for 2022 at 52 $\rm CO_{2}\text{-e}/kWh$}, the added impact is about 104 kg of $\rm CO_{2}\text{-e}$}.
@@ 6 Conclusion
In this paper, we generalized our YOLO-CIANNA deep learning and characterization method to 3D hyperspectral data from the SKAO SDC2. We detail the required modifications to the box definition, the position and size regression loss subpart, and the match association metric. We present a backbone optimized for HI line emission detection and describe the construction of our training sample, along with our training and inference pipelines.
Our results show that 3D CNN structures efficiently extract complex 3D patterns specific to HI sources and are capable of best-in-class source characterization. However, we highlight that implementing such a method on complex astronomical data is difficult and that only a narrow region of the parameter and architecture space allows successful training. We also introduce a bootstrap training strategy that relies on the self-assessed detection capability of a first poorly constrained model to refine the training selection function iteratively.
These state-of-the-art results are encouraging regarding the generalization of the method to observed surveys from SKA precursor instruments, with preparatory work currently underway on the LADUMA and WALLABY surveys. Although 3D CNN approaches should not be considered drop-in replacements for classical analysis tools, which often offer greater interpretability and consistency, our results demonstrate that deep learning detectors can efficiently produce complete and pure first-guest catalogs from large hyperspectral datasets.
Example scripts and notebooks are provided in the CIANNA git repository to help reproduce our results. We also publish all detection catalogs and trained models presented in this paper at xx.xxxx/zenodo.xxxxxxxx [^4] Pre-review archive at https://share.obspm.fr/s/swyCT7BgEGjtZK3, along with ancillary models trained over the whole MAIN cube to serve as pre-training states for transfer learning toward other data or surveys (Appendix E).
@@ Acknowledgements.
We acknowledge the support granted to the MINERVA project by the Paris Observatory, including material and financial support. We acknowledge the Paris Observatory DIO shared computing resources and their handling of the MINERVA dedicated compute resources. We acknowledge the computing resources from the GENCI (Jean-Zay GPU partition) provided to the MINERVA team in the context of the SDC2. DC acknowledges the MINERVA project for funding a two-year post-doctoral contract. DC acknowledges its PSL fellowship as part of the MCA-IA program and in collaboration with PR[AI]RIE-PSAI. We highlight the great work done by SKAO in organizing all the SDCs and thank them for allowing post-challenge access to the datasets.
@@ Appendix A Source parameter space correlations
Both the MAIN and LDEV truth catalogs exhibit correlations between source parameters. As we observed in Sect. 4.2, the inclination angle error distribution exhibits substructures, indicating that the network likely infers this parameter from other source properties. We identified parameter pairs showing specific correlations and represent them in the form of 2D histograms for all matched sources from our BT1 catalog in Fig 7. To determine if our detector reconstructs these correlations, we show the distribution for both the predicted and associated target values. We also display histograms for candidate detections rejected by the scorer’s matching criteria and for the training sample to determine if the parameter space coverage allows for the identification of these correlations.
We observe a linear correlation between the flux and the HI source size. We also notice substructured distributions in both histograms that involve the inclination angle. Our detector fails to capture most fine-grained substructures, reducing them to mostly linear relations, which is likely due to the sparse sampling of the parameter spaces in the training sample. Still, the observed relations confirm that the model learns to correlate source properties, allowing it to infer some parameters from others that are easier to evaluate. Finally, we observe that most false detections are faint sources with a small HI size and a line width that falls in the lower half of the distribution. Although it is not illustrated here, we note that false detections are uniformly distributed in sky position and central frequency.
Figure 7: Two-dimensional histograms of selected source parameter pairs. Rows correspond to different source catalogs: first and second rows represent matched sources from the MAIN from BT1 using the predicted and target source values, respectively; third row represents false detections; fourth row represents the selected training catalog from the LDEV cube.
@@ Appendix B Local reprojection effect on detectability
Figure 8: Signal projection for a subset of matched MAIN-cube detections. Each row corresponds to a source. The left column is obtained from cutouts centered on the source in the original MAIN cube, whereas the right column is obtained after reprojecting the cutout based on each source’s PAPA and $i$. For each cutout, we present two 2D projections, one in the plane of the sky by averaging over the frequency axis, and the second in a sky-frequency plane by averaging over one sky axis. Images are in arbitrary units inversely proportional to the normalized voxel intensity. The colormap is adjusted to maximise the per-image contrast. The truth catalog id and the target source properties are indicated for reference. Boxes are overplotted on the left column for each source and represent the target box in green and the corresponding detector’s prediction in blue.
@@
Most participating teams started by visually exploring the SDC2 cubes, identifying only a handful of very bright sources. Smoothing the LDEV in the frequency axis can increase the number of visually identified sources up to a hundred ($\sim 1\%$ of the LDEV catalog). Visualizing the source signal in 3D, like we did with Fig. 1, is only possible for the brightest sources. Fainter sources can be visualized by taking a cutout centered on the source coordinate and projecting their signal along one of the cube axes based on the source extension, which boosts the S/N. This per-source approach obviously outperforms global cube smoothing in enhancing the individual source signal. However, it requires knowing the position and extension of each source in advance, defeating the very purpose of a detection and characterization task. Still, this test confirms that around 10% of sources technically contribute enough signal to be detectable, which is still lower than our detector capabilities.
@@
Here, we aim to identify potential explanations for the robust detection capability of our method on this task. We hypothesize that 3D CNNs can learn to recombine the cube information in a way that allows them to explore multiple 3D reprojections and various source extensions dynamically based on the input context. To test this, we can try to visualize reprojected sources into their preferred plane based on the truth angles. We do this for a few successfully detected sources from the MAIN cube and present the results in Fig. 8. To better illustrate the effect of the reprojection, we select sources covering an extensive range of S/N and compare their 2D projections in the original cube and in the projected cutout. The first source is the same as the one represented in Fig. 1. Horizontal periodic lines visible in the reprojected space are artifacts resulting from the reprojection method we used. Boxes corresponding to our target definition and the associated network prediction are overplotted on the projections from the original cube.
@@
As expected, we see that reprojecting based on known source characteristics boosts the signal. Any classical detection methods applied directly to this reprojected space would likely achieve similar performance to our detector. The strength of our method lies in its ability to identify such transformations automatically on the fly for each source. The fully convolutional 3D translational-equivariant structure of our backbone implies that our detector behaves similarly to simultaneously applying a given detector centered on every subregion mapped by our network, as defined by the output grid. Considering that this detector is built upon stacks of 3D convolutions, it can simultaneously scan for a vast diversity of 3D patterns that could typically emulate the expected projections. Achieving the same result with a classical method would require following a sliding window approach, testing all possible positions in the 3D cubes with multiple detector configurations to account for the possible reprojections, which would be technically prohibitive in terms of compute time.
@@ Appendix C Detection-only task performances
In Sect. 4.2, we discussed how the SDC2 scorer correlates the detection and source characterization and how it affects the evaluation of the detector’s capability. Here, we test whether merging detection and characterization into a single model compromises pure detection performance. This question arises because the YOLO-CIANNA loss combines multiple objectives balanced by scaling factors (Eq. 14 in Paper I), which may compete against each other. This question is addressed for the specific case of the SDC1 in Appendix D of Paper I, where a DIoU-based AP score is used to evaluate the purely geometrical detection capabilities of our method across different training setups. It concludes that our combined detection and characterization model performs systematically better than equivalent models for which the parameter prediction subloss is deactivated. Although a naïve combined loss would lead to slightly lower detection performances, the cascading loss design in YOLO-CIANNA leverages the additional information, improving both detection and characterization results. These observations align with our general hypothesis that simultaneously predicting correlated variables often yields better results than making independent predictions.
In this section, we do the same verification for the SDC2 by computing purely geometrical 3D DIoU-based AP scores at different thresholds for our three reference models from Sect. 4.1, to which we add a model trained with parameter prediction turned off (No-param). Our bootstrap training approach is based on matches provided by the scorer, which requires source parameter predictions. Consequently, this new model is trained using the BASE model configuration and with our default selection function (Sect.  2.3). We compute the AP scores for DIoU thresholds of -0.3, 0.1, and 0.5, which correspond to the DIoU limits used in the second and first NMS of the prediction pipeline, and to a strict high-quality match criteria, respectively. The absence of source characteristic predictions prevents us from using the scorer to identify the best iteration. We therefore rely directly on the AP scores for the No-param model. Since we use the previously identified best iteration for all other models, this might grant a slight unfair advantage to the No-param scores. We present the AP scores in Table 5, along with the scorer-based AP for our three reference models. The model without source characterization matches our BASE model AP scores for the first two thresholds but falls behind for the stricter one. It appears that the added parameter prediction helps improve source positioning and sizing, but has no substantial impact on source detectability. Still, we can confirm that simultaneous detection and characterization have no adverse effect on pure detection accuracy using our YOLO-CIANNA method on the SDC2.
@@
Appendix D Bootstrap from a degraded selection function
We show in Sect. 4.1 that our bootstrap training approach introduced in Sect. 3.4 allows us to improve the score from our BASE model significantly. However, the improvement may appear modest with respect to the increased training time and pipeline complexity. We hypothesize that our initial selection function is already near optimal, and that the bootstrap method would allow us to converge toward a similar final score if we started from a poorer selection. To test this, we define a degraded selection function as

### v2

@@ 5 Discussion
This second edition of the SKAO SDCs improves the challenge format in several ways compared to the SDC1. The simulated cube includes more physical and instrumental effects, improving its realism, which is also a strength of the subsequent SDC editions (see the SDC3a, Bonaldi et al., 2025). The LDEV cube is a proper representation of the MAIN cube, allowing one to train supervised ML methods without requiring representativity corrections like we implemented in Paper I. Still, the use of a similar scoring method to the SDC1 results in the same bias and limits, such as strong coupling between detection and characterization performance evaluation, or the absence of a confidence score weighting like in most computer-vision challenges. We also acknowledge that using simulated data provides an unrealistic advantage for training supervised ML methods compared to real observations. All sources are known and perfectly characterized, allowing for the definition of unambiguous targets, which would not be the case with observational labeled catalogs.
@@
The results obtained with our YOLO-CIANNA 3D method, combined with the fact that the two top-scoring teams from the SDC2 utilized 3D CNNs, demonstrate that this type of structure offers significant advantages for processing noisy hyperspectral data. However, as discussed in Paper I, this requires one to build dedicated approaches that account for the specificities of astronomical data at the heart of their design. For this study, this was facilitated by the low-level nature of our CIANNA framework. Nevertheless, several aspects of our method could be improved. In terms of backbone architecture, CIANNA V-1.0 does not implement residual and skip connection layers (He et al., 2016), which limit the maximum network depth. Such layers have been added to the experimental build of CIANNA, which should be publicly released on a short timescale. Unlike some versions of YOLO (Redmon and Farhadi, 2018), we did not split predictions across scales, which is also related to current architectural limitations. In addition, having a fixed number of detection units on a finite-sized grid constrains both the positioning resolution and the density of detectable sources. This limit could likely be overcome by adding attention layers (Vaswani et al., 2017; Carion et al., 2020; Zhang et al., 2021) and modifying the definition of the output layer. We are also exploring pre-training of our backbone using the denoising diffusion probabilistic model (DDPM, Ho et al., 2020) formalism. This could be used to pre-train a model over the MAIN cube without requiring any source targets. Finally, we could inject context information, such as the instrumental setup, at multiple stages of the backbone to guide the source extraction (Lu et al., 2019). This would enable us to train generic detection models compatible with multiple setups or instruments.
@@
The natural next step after achieving state-of-the-art results on simulated data is to generalize our method application to observational data from SKA precursors such as the LADUMA (Looking at the Distant Universe with the MeerKAT Array, Blyth et al., 2016) and WALLABY (Widefield ASKAP L-band Legacy All-sky Blind surveY, Koribalski et al., 2020) surveys. Our method could also be generalized to detect line emission in hyperspectral surveys covering a different frequency domain, such as MUSE. However, it requires a large set of labeled examples to build the training and validation sets. Although simulations can generate large example datasets for pre-training highly parametric statistical models, they often lose accuracy when generalized to real observations. A complementary training phase on real labeled data is almost always necessary, and the inference pipeline must necessarily be calibrated on a labeled real validation dataset. For this, we would combine labels from classical detection methods, crowdsourced science, or existing catalogs from other instruments. The supervised nature of our method implies that the trained detector would likely inherit uncertainties and biases from the methods used to define the targets. Combining independent labeling methods might help in this regard. In all cases, the bootstrap training strategy can be used to train a model from an imperfect training sample to identify new candidates, which can then be confirmed using other labeling approaches or new observational programs, iteratively improving detection sensitivity. Preparatory work indicates that direct porting of our SDC2-trained models generalizes well to preliminary LADUMA data after a simple rescaling, producing candidate detections compatible with catalogs obtained with SOFIA. Specific retraining for this survey is currently being explored.
[…]
@@ 6 Conclusion
In this paper, we generalize our YOLO-CIANNA deep learning and characterization method to 3D hyperspectral data from the SKAO SDC2. We detail the required modifications to the box definition, the position and size regression loss subparts, and the match association metric. We present a backbone optimized for HI line-emission detection and describe the construction of our training sample and our training and inference pipelines.
Our results show that 3D CNN structures efficiently extract complex 3D patterns specific to HI sources and achieve best-in-class source characterization. However, we highlight that implementing such a method on complex astronomical data is challenging in practice, and that only a narrow region of the parameter and architecture space enables successful training. We also introduce a bootstrap training strategy that relies on the self-assessed detection capability of successively trained models to refine the selection function used to build the training sample.
These state-of-the-art results are encouraging regarding the generalization of the method to observed surveys from SKA precursor instruments, with preparatory work currently underway for the LADUMA and WALLABY surveys. Although 3D CNN approaches should not be considered drop-in replacements for classical analysis tools, which often offer greater interpretability and consistency, our results demonstrate that deep learning detectors can efficiently produce complete and pure first-guest catalogs from large hyperspectral datasets.
7 Data availability
Example scripts and notebooks are provided in the CIANNA git repository to help reproduce our results. We also publish all training and inference codes, trained models, and detection catalogs presented in this paper at 10.5281/zenodo.18403011, along with ancillary models trained over the whole MAIN cube to serve as pre-training states for transfer learning toward other data or surveys (Appendix F).
@@ Acknowledgements.
We acknowledge the support granted to the MINERVA project by the Paris Observatory, including material and financial support. We acknowledge the Paris Observatory DIO shared computing resources and their handling of the MINERVA dedicated compute resources. We acknowledge the computing resources from the GENCI (Jean-Zay GPU partition) provided to the MINERVA team in the context of the SDC2. DC acknowledges the MINERVA project for funding a two-year post-doctoral contract. DC acknowledges its PSL fellowship as part of the MCA-IA program and in collaboration with PR[AI]RIE-PSAI. We highlight the great work SKAO has done in organizing all the SDCs and thank them for allowing post-challenge access to the datasets.
@@ Appendix A Source parameter space correlations
Both the MAIN and LDEV truth catalogs exhibit correlations between source parameters. As we observed in Sect. 4.2, the inclination angle error distribution exhibits substructures, suggesting that the network likely infers this parameter from other source properties. We identified parameter pairs that exhibit correlations and present them in the form of 2D histograms for all matched sources in our BT1 catalog in Fig. 7. To determine whether our detector reconstructs these correlations, we show the distributions for both the predicted and associated target values. We also display histograms for candidate detections rejected by the scorer’s matching criteria and for the training sample to determine if the parameter space coverage is sufficient for these correlations to be clearly identified.
We observe a linear correlation between the flux and the HI source size. We also observe substructured distributions in both histograms that involve the inclination angle. Our detector fails to capture most fine-grained substructures, reducing them to mostly linear relations, which is likely due to the sparse sampling of the parameter space by the training sample. Still, the observed relations confirm that the model learns to correlate source properties, allowing it to infer some parameters from others that are easier to evaluate. Finally, we observe that most false detections are faint sources with a small HI size and line width that fall in the lower half of the distribution. Although not shown here, we note that false detections are uniformly distributed across sky position and central frequency.
Figure 7: Two-dimensional histograms of selected source parameter pairs. Rows correspond to different source catalogs: the first and second rows represent matched sources from the MAIN from BT1 using the predicted and target source values, respectively; the third row represents false detections; the fourth row represents the selected training catalog from the LDEV cube.
@@ Appendix B Local reprojection effect on detectability
Figure 8: Signal projection for a subset of matched MAIN-cube detections. Each row corresponds to a source. The left column is obtained from cutouts centered on the source in the original MAIN cube, whereas the right column is obtained after reprojecting the cutout based on each source’s PAPA and $i$. For each cutout, we present two 2D projections, one in the plane of the sky by averaging over the frequency axis, and the second in a sky-frequency plane by averaging over one sky axis. Images are in arbitrary units inversely proportional to the normalized voxel intensity. The colormap is adjusted to maximize the per-image contrast. The truth catalog id and the target source properties are indicated for reference. Boxes are overplotted on the left column for each source and represent the target box in green and the corresponding detector’s prediction in blue.
@@
Like most participating teams, upon getting access to the SDC2 data, our first action was to explore the cubes visually. This approach allowed us to identify only a handful of very bright sources. Smoothing the LDEV cube in the frequency axis can increase the number of visually identified sources up to a hundred ($\sim 1\%$ of the LDEV catalog). Directly visualizing the source signal in 3D, as we did with Fig. 1, is only possible for the brightest sources. Fainter sources can be visualized by taking a cutout centered on the source coordinate and projecting their signal along one of the cube axes based on the source extension, which boosts the S/N. This per-source approach obviously outperforms global cube smoothing in enhancing the individual source signal. However, it requires prior knowledge of each source’s position and extension, which defeats the purpose of a detection and characterization task. Still, this test confirms that around 10% of sources have sufficiently high signal density to be detectable, which remains lower than the capabilities of our YOLO-CIANNA detector.
@@
Here, we aim to identify potential explanations for the robust detection capability of our method on this task. We hypothesize that 3D CNNs can learn to recombine the input signal in a way that enables them to dynamically explore multiple 3D reprojections and various source extensions based on the input context. To test this, we can identify the source main signal plane using the known truth PA and inclination angles, and reproject the cube to align with it. We do this for a few successfully detected sources from the MAIN cube and present the results in Fig. 8. To better illustrate the effect of this per-source reprojection, we select sources covering an extensive range of S/N and compare their 2D projections in the original cube and in the reprojected cutouts. The first source is the same as the one represented in Fig. 1. Regular horizontal lines visible in the reprojected space are artifacts of the cube reprojection method we used. Boxes corresponding to our target definition and the associated network prediction are overplotted on the projections from the original cube.
@@
As expected, reprojecting based on known source characteristics boosts the signal. Any classical detection methods applied directly to this reprojected space would likely achieve similar performance to that of our detector. The strength of our method lies in its ability to identify such transformations automatically on the fly for each source. The fully convolutional 3D translational-equivariant structure of our backbone implies that our detector behaves as if the same sub-detector were simultaneously applied to every subregion mapped by our network, as defined by the output grid. Because this sub-detector is built from stacks of 3D convolutions, it can simultaneously scan for a wide range of 3D patterns that can typically emulate the expected projections. Achieving the same result with a classical method would require one to follow a sliding window approach, testing all possible positions in the 3D cubes and multiple detector configurations to account for the variety of suitable reprojections, which would be computationally prohibitive.
@@ Appendix C Detection-only task performances
In Sect. 4.2, we discussed how the SDC2 scorer correlates the detection and source characterization and how it affects the evaluation of the detector’s capability. Here, we evaluate whether merging detection and characterization into a single model compromises pure detection performance. This question arises because the YOLO-CIANNA loss combines multiple objectives that are balanced through scaling factors (Eq. 14 in Paper I), which may compete against each other. This question is addressed for the specific case of the SDC1 in Appendix D of Paper I, where a DIoU-based AP score is used to evaluate the purely geometric detection capabilities of our method across different training setups. It concludes that our combined detection and characterization model performs systematically better than equivalent models for which the parameter prediction subloss is deactivated. Although a naive combined loss could slightly reduce detection performance, the cascading loss design in YOLO-CIANNA ensures that the characterization target is used only for sources that are already confidently detected, thereby improving both detection and characterization results. These observations align with our general hypothesis that simultaneously predicting correlated variables often yields better results than making independent predictions.
In this section, we present the results of a similar verification for the SDC2, computed using purely geometric 3D DIoU-based AP scores at different thresholds for our four reference models from Sect. 4.1, to which we added a model trained with parameter prediction turned off (No-param). This new model was trained using the BASE model configuration and with our default selection function (Sect. 2.3). The absence of parameter predictions prevents bootstrapping, as it relies on the scorer matching criteria. We compute AP scores for DIoU thresholds of -0.3, 0.1, and 0.5, which correspond to the DIoU limits used in the second and first NMS steps of the prediction pipeline, and to a strict high-quality match criterion, respectively. The absence of source characteristic predictions also prevents us from using the scorer to identify the best iteration. We therefore rely directly on the geometric AP scores to select the best iteration for the No-param model, while keeping the best iteration based on the SDC2 scorer for all other models. We note that this choice might give the No-param model a small unfair advantage regarding AP scores. We present the purely geometric AP scores in Table 6, along with the scorer-based AP for our four reference models. The model without source characterization matches our BASE model’s geometric AP scores at the first two thresholds but falls behind for the stricter threshold. It appears that the added parameter prediction improves source positioning and sizing but has no substantial impact on source detectability. Still, we can confirm that simultaneous detection and characterization have no adverse effect on pure detection accuracy using our YOLO-CIANNA method on the SDC2.
@@
Appendix D Bootstrap with from-scratch training
In Sect. 3.5, we stated that during bootstrap, we reuse part of the trained layers from the previous model at each step, simply reinitializing the last three layers with random weights. This approach allows us to start from a model already capable of identifying obvious objects, removing the initial training plateau observed in the BASE model’s validation loss in Fig. 4. The training can then focus on identifying fainter objects added to the training sample based on previous model detections. Here, we provide a reference point for what happens when starting with a fully randomized set of weights for all layers (from-scratch training) during bootstrapping. To exacerbate the effects, we retrained the equivalent of our BT2 model, which corresponds to a second bootstrap step, creating an FS-BT2 model. We represent the evolution of the validation loss as a function of the training iteration in Fig. 9.
We observe an initial loss plateau for around 800 iterations, indicating that the model has difficulties in understanding the task and detecting even the brightest objects. This is mostly due to the addition of fainter targets to the training sample, which reduces the proportion of cases with an obvious object and adds training noise by requiring the model to detect sources in regions dominated by noise. This is typically the regime that the starting selection function tries to avoid. While flagging added sources as difficult helps mitigate the effect, it still affects training, especially given the reduction in the proportion of easy cases presented to the model due to our training example generation strategy (see Sect. 3.4). We also observe that the maximum score achieved by FS-BT2 (24 653 points with 33 228 matches and 2606 false detections) is measurably lower than that of our BT2 model and closer to that of our BASE model, and that it is obtained at a later iteration than BT2. In theory, as long as the model can break out of the initial plateau, it should converge to a score similar to that of BT2, since they share the same training sample. We note that the average source score is lower than that of all other models, with $\bar{s}=0.8201$, but this accounts for only half of the score difference ($\approx$300 points). The rest most likely comes from a specific dynamic of the training: some YOLO-CIANNA loss subparts likely start overtraining (likely the parameter prediction) before others converge, due to the delayed beginning of global convergence. Such behavior could be addressed by adjusting the loss subpart scaling or other hyperparameters, but this would create an inconsistency in the training setup across successive bootstrap steps. From these observations, we consider our default approach of reusing part of the previously trained model more efficient than systematic from-scratch retraining.
Figure 9: Evolution of the validation loss subparts (natural, Sect. A.7 in Paper I) during training for our FS-BT2 model. The first 20 iterations are omitted. Scores computed over the MAIN cube are indicated every 200 iterations starting at 1000.
Appendix E Bootstrap from a degraded selection function
We show in Sect. 4.1 that our bootstrap training approach introduced in Sect. 3.5 significantly improves the score of our BASE model. However, the improvement may appear modest given the increased training time and pipeline complexity. We hypothesize that our initial selection function is already near optimal, and that the bootstrap method would allow us to achieve a similar final score if we started from a poorer selection. To test this, we defined a degraded selection function as